OpenStudy (anonymous):

A spherical snowball is melting in such a way that its diameter is decreasing at rate of 0.4 cm/min. At what rate is the volume of the snowball decreasing when the diameter is 12 cm? Rate of change of volume = cm3/min

OpenStudy (anonymous):

im confused about what the formula is for volume

OpenStudy (anonymous):

@jim_thompson5910 can you help me out here?

OpenStudy (anonymous):

?????? why u dont use the formula for the volume of a sphere?

OpenStudy (anonymous):

so V = (4 pi a^3)/3

OpenStudy (anonymous):

yes and u must get the derivative with respetc to time i meant dV/dt

OpenStudy (anonymous):

\[\frac{ dV }{ dt }=(4pir^2)\frac{ dr }{ dt }\]

OpenStudy (anonymous):

but i need this in diameter

OpenStudy (anonymous):

@julian25

OpenStudy (anonymous):

hey Zepdrix

zepdrix (zepdrix):

Hey c:

OpenStudy (anonymous):

how have you been doing?

zepdrix (zepdrix):

\[\large V=\frac{4}{3}\pi r^3\] They told us that,\[\large d'=-0.4\] Understand why this is negative?

zepdrix (zepdrix):

Good good c:

OpenStudy (anonymous):

because its decreasing

zepdrix (zepdrix):

So they've told us to find \(\large V'\) when \(\large d=12\)

zepdrix (zepdrix):

So we just need to rewrite our volume formula in terms of diameter right?

OpenStudy (anonymous):

ya

zepdrix (zepdrix):

How does diameter relate to radius? :D

OpenStudy (anonymous):

d=2r d d = 2 d r

OpenStudy (anonymous):

d=2r

zepdrix (zepdrix):

Hmm ok let's try using that, don't forget about the `cube` being applied to the entire substitution. \[\large V=\frac{4}{3}\pi r^3 \qquad \rightarrow \qquad V=\frac{4}{3}\pi \left(\frac{1}{2}d\right)^3\]

OpenStudy (anonymous):

oops its actually d d/dt = 2 dr/dt

zepdrix (zepdrix):

I skipped a step in there, I hope that wasn't confusing.\[\large d=2r \qquad \rightarrow \qquad \frac{1}{2}d=r\]And then I made a substitution for the r.

OpenStudy (anonymous):

na your good

OpenStudy (anonymous):

V=4/3π(1/2d)^3 so dV/dt=4pi(1/2d)^2

zepdrix (zepdrix):

So you should be able to simplify that down a bit.. I think it gives us something like this,\[\large V=\frac{1}{6}\pi d^3\] I think we want to simplify before we take a derivative. Otherwise you'll need to remember to apply the chain rule. If you're comfortable with that, then it's fine.

zepdrix (zepdrix):

Cause I can see in your dV/dt, you didn't apply the chain rule. A factor of 1/2 should be popping out from that 1/2d due to the chain rule.

OpenStudy (anonymous):

ok, ya i suck at this stuff

OpenStudy (anonymous):

so we would have 1/2pi(d)^2

zepdrix (zepdrix):

Either way works fine. If we do it the way you did, we'll apply the chain rule to the inside.\[\large \frac{dV}{dt}=4\pi\left(\frac{1}{2}d\right)^2\frac{d}{dt}\left(\frac{1}{2}d\right)\]

zepdrix (zepdrix):

Looks good, you're missing one really important piece of information though.

zepdrix (zepdrix):

we should have called our diameter D, I'm realizing that now. lol

OpenStudy (anonymous):

never to late lol

zepdrix (zepdrix):

What happens when you take the derivative of diameter with respect to `time`. The same thing that happens when you take the derivative of Y with respect to X. Since it's in respect to a different variable, we need to multiply it by the dD/dt term.

zepdrix (zepdrix):

I really really prefer prime notation, it makes it so much easier to read. Let's try that real quick. \[\large V=\frac{4}{3}\pi \left(\frac{1}{2}d\right)^3\]Taking the derivative we found to be,\[\large V'=4\pi\left(\frac{1}{2}d\right)^2\left(\frac{1}{2}d\right)'\]Which gives us,\[\large V'=2\pi\left(\frac{1}{2}d\right)^2d'\]

OpenStudy (anonymous):

ok, but now where does -0.4 cm/min go?

zepdrix (zepdrix):

Given: \(\large d=12\) \(\large d'=-0.4\) Solve for \(\large V'\) Just plug and chug!

OpenStudy (anonymous):

ah ok got it

zepdrix (zepdrix):

Ya you got it! c: too quick for me hah

OpenStudy (anonymous):

so dV/dt=2π(1/2(12))^2(-.4)

zepdrix (zepdrix):

Looks good.

OpenStudy (anonymous):

ok tyvm, you got time for another?

OpenStudy (anonymous):

?

zepdrix (zepdrix):

Oh sorry was microwaving some soup c: Sure